OpenStudy (ellableuuu):

Write a function rule based on the table below. x f(x) 0 0 2 4 4 16 6 36 - Select one: a. f(x) = x2 b. f(x) = 2x c. f(x) = x + 2 d. f(x) = 4x - My answer is A, is that correct?

OpenStudy (sooobored):

only if answer a is f(x)= x^2 \(=x^2\)

OpenStudy (ellableuuu):

It is, so am I right?

OpenStudy (ellableuuu):

@sooobored

OpenStudy (ellableuuu):

@pooja195

OpenStudy (mathmale):

Well, yes and no. f(x) = x2 needs to be re-written as f(x) = x^2. We need to use the " ^ " symbol to denote exponentiation. This is important to avoid ambiguity.

OpenStudy (mathmale):

But yes, your answer choice is correct.